The demand function for bicycles in Holland has been estimated to be Q = 2,000 + 15Y – 5.5P where Y in income in thousands of euros, Q is the quantity demanded in units, and P is the price per unit. When P = 150 euros and Y = 15(000) euros, determine the following:

a. Price elasticity of demand

b. Income elasticity of demand

 



1
Expert's answer
2021-03-09T15:17:59-0500

Let's first find the quantity demanded at "P=150" and "Y=15":


"Q=2000+15\cdot15-5.5\cdot150=1400."

(a) The price elasticity of demand can be found as follows:


"\epsilon=\dfrac{P}{Q}\cdot\dfrac{dP}{dQ},""\epsilon=\dfrac{150}{1400}\cdot(-5.5)=-0.58."

(b) The income elasticity of demand can be found as follows:


"\epsilon=\dfrac{I}{Q}\cdot\dfrac{dQ}{dI},""\epsilon=\dfrac{15}{1400}\cdot15=0.16."